A cubical block of mass 2.5kg is kept on a levelled horizontal surface . If exerts a pressure of 100pa on the surface then find the volume of the block
1
Expert's answer
2020-07-23T08:46:08-0400

Let's first find the area of the cubical surface from the definition of the pressure:


"p = \dfrac{F}{A},""A = \dfrac{F}{p} = \dfrac{mg}{p} = \dfrac{2.5 \ kg \cdot 10 \ \dfrac{m}{s^2}}{100 \ Pa} = 0.25 m^2."

Then, we can find the side length of the cubical block:


"a = \sqrt{A} = \sqrt{0.25 \ m^2} = 0.5 \ m."

Finally, we can calculate the volume of the block:


"V = a^3 = (0.5 \ m)^3 = 0.125 \ m^3."

Answer:

"V = 0.125 \ m^3."